OpenStudy (anonymous):

Solve and see whether or not it is extraneous 4/x^2+3x-10 -1/x^2+x-6 = 3/x^2-x-12

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

now it gets real annoying real fast factor the denominators first i suck at factoring so let me know what you get

OpenStudy (anonymous):

Alright hold on lol:)

OpenStudy (anonymous):

4/(x-2)(x+5) -1/(x-2)(x+3)= 3(x-4)(x+3) @satellite73 done :)

OpenStudy (anonymous):

whew ok

OpenStudy (anonymous):

then you can clear the fractions as before but this time you have to multiply by the least common multiple of the denominators which is \[(x-2)(x+5)(x+3)(x-4)\] ugh

OpenStudy (anonymous):

oh..oh gosh...headachee

OpenStudy (anonymous):

actually it is not that bad, since the numerators are all numbers

OpenStudy (anonymous):

\[4(x+3)(x-4)-1(x+5)(x-4)=3(x+5)(x-2)\] i think check it

OpenStudy (anonymous):

ok hold on:)

OpenStudy (anonymous):

it is really just cancelling what you need to cancel

OpenStudy (anonymous):

yes, thats true

OpenStudy (anonymous):

then we have a raft of algebra to do multiply all this muck out, combine like terms etc etc i refuse to do it, i would cheat

OpenStudy (anonymous):

Haha! I totally agree:) Ok, lets move onto another one

OpenStudy (anonymous):

lol actually all the \(x^2\) terms go bye bye and you end up with only \(14x=2\) so \(x=\frac{1}{7}\)

OpenStudy (anonymous):

and since \(\frac{1}{7}\) does not make any denominator zero, it is NOT extraneous

OpenStudy (anonymous):

Marry me? Lol jk! Thanks so much :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

and sure why not? it can be an OS wedding

OpenStudy (anonymous):

Perfect! I'll start planning :D